write an equation of the line containing the given point and parallel to the given line. Express your answer in the form y=mx+b (8,9); x+2y=7
Answer by mananth(16946)

1 x + 2 y = 7
Find the slope of this line
make y the subject
2 y = -1 x + 7
Divide by 2
y = -0.5 x + 3.5
Compare this equation with y=mx+b
slope m = -0.5 -0.5
The slope of a line parallel to the above line will be the same
The slope of the required line will be -0.5
m= -0.5 ,point ( 8 , 9 )
Find b by plugging the values of m & the point in
y=mx+b
9 = -4.00 + b
b= 13
m= -0.5
Plug value of the slope and b in y = mx +b
The required equation isy = -0.5 x + 13
y = - 1/ 2 x + 13
m.ananth@hotmail.ca